# Attractor learning for spatiotemporally chaotic dynamical systems using echo state networks with transfer learning

### v1

@@ 4.2.2 Predicting the power spectrum of the KS equation using transfer learning
Figure 2: Transfer learning allows the ESN to accurately predict the power spectrum of the KS equation as the size of the spatial domain varies. Curves illustrate power spectra. DNS-L43: Power spectrum obtained by averaging over the $10$ trajectories in the $L=43$ regime used for prediction. ESN-L22: Power spectrum for the $L=22$ regime predicted by the ESN, once it is trained on $20$ trajectories from the $L=22$ regime. ESN-TL: Power spectrum for the $L=43$ regime predicted by the ESN after transfer learning of level $10\%$ (left), $25\%$ (middle), and $50\%$ (right). Plots show $\log(e_{k})$ versus $\log(k)$, where $k$ denotes wavenumber.
4.3 Predicting the power spectrum of the gKS equation
We consider the gKS equation when the parameter $\gamma$ in the dispersion relation is small. We assume that for $\gamma$ small, the gKS dynamics admit an attractor that supports an ergodic invariant measure. We assume that this invariant measure describes the asymptotic distribution of all of the initial data that we select. We show that for $\gamma$ small and fixed, the ESN accurately predicts the power spectrum of the gKS equation for this fixed value of $\gamma$, once trained on trajectories generated using this fixed value of $\gamma$. We then show that transfer learning is remarkably effective for small $\gamma$. Throughout Section 4.3, we set $L=43$.
4.3.1 Predicting the power spectrum of the gKS equation without transfer learning
We performed the following experiment for each of two values of $\gamma$, $\gamma=0$ and $\gamma=0.1$. We began with an untrained ESN. We generated $30$ gKS trajectories, each of length $T=1000$ Lyapunov times, via direct numerical simulation. We used $20$ of these to train the ESN and the other $10$ for prediction.
Figure 3 illustrates that for $\gamma=0$ (left) and $\gamma=0.1$ (right), the trained ESN accurately predicts the power spectrum of the gKS equation. Each red curve shows the ESN prediction of the power spectrum. Each blue curve shows the power spectrum obtained by averaging over the $10$ trajectories used for prediction. Importantly, the power spectrum of the gKS equation depends sensitively on $\gamma$ in the chaotic (small $\gamma$) regime. This observation serves as serious motivation for the use of transfer learning in this regime.
@@ 4.3.1 Predicting the power spectrum of the gKS equation without transfer learning
Figure 3: Once trained, the ESN accurately predicts the power spectrum of the gKS equation. Left: $\gamma=0$. Right: $\gamma=0.1$. Each blue curve shows the power spectrum obtained by averaging over the $10$ DNS trajectories used for prediction. Each red curve shows the power spectrum predicted by the ESN, after the untrained ESN has been trained using $20$ trajectories. Plots show $\log(e_{k})$ versus $\log(k)$, where $k$ denotes wavenumber.
4.3.2 Predicting the power spectrum of the gKS equation with transfer learning
Figure 3 demonstrates that the statistical behavior of the gKS equation depends sensitively on $\gamma$ in the small-$\gamma$ regime. We therefore ask if transfer learning can capture substantial changes in statistical behavior as $\gamma$ varies. To address this question, we performed the following experiment. We started with an untrained ESN and first trained it using $20$ trajectories from the $\gamma=0$ regime. We then updated the training using a small amount of TL data from the $\gamma=0.1$ regime at transfer learning rate $\alpha=0.005$.
@@ 4.3.2 Predicting the power spectrum of the gKS equation with transfer learning
Figure 4: Transfer learning allows the ESN to accurately predict the power spectrum of the gKS equation as the dispersion parameter $\gamma$ varies. Blue curve (same curve in both panels): Power spectrum of the gKS equation in the $\gamma=0.1$ regime obtained by averaging over the $10$ DNS trajectories used for prediction. Green curve (same curve in both panels): Power spectrum predicted by the ESN after a first round of training using $20$ trajectories from the $\gamma=0$ regime and then a training update using TL data from the $\gamma=0.1$ regime at level $10\%$. Yellow curve: Power spectrum predicted by the ESN after training only on $20$ trajectories from the $\gamma=0$ regime. Red curve: Power spectrum predicted by an ESN that has been trained only on the TL data, namely $2$ trajectories from the $\gamma=0.1$ regime. Plots show $\log(e_{k})$ versus $\log(k)$, where $k$ denotes wavenumber.
@@
The result of this experiment is shown in Figure 4 (left). The blue curve shows the ground truth, namely the power spectrum in the $\gamma=0.1$ regime obtained by averaging over the $10$ trajectories from the $\gamma=0.1$ regime used for prediction. Using a transfer learning level of only $10\%$ ($2$ trajectories) produces an accurate ESN prediction of the power spectrum (green curve). For reference, the yellow curve shows the ESN-based prediction of the power spectrum when the ESN is trained only on $20$ trajectories from the $\gamma=0$ regime. Importantly, the accurate TL-based prediction of the power spectrum is valid over long timescales. By contrast, single-orbit predictions in the $\gamma=0.1$ regime based on transfer learning are valid for shorter timescales [1, Section 5.1.1].
@@
As a control, we compare the prediction using transfer learning (Figure 4, green curve, ESN-TL) to a prediction from an ESN that has been trained only on the TL data, namely $2$ trajectories from the $\gamma=0.1$ regime (Figure 4, red curve, ESN*). We see that the ESN* prediction is inferior to the ESN-TL prediction in the eye metric. Discrepancies between these two predictions occur at low wavenumbers that contain most of the energy. Quantitatively, the curves shown in Figure 4 (right) contain total energies $2.92$ (blue, DNS), $2.80$ (red, ESN*), and $2.97$ (green, ESN-TL). This corresponds to relative error $4.1\%$ and $1.7\%$ for the ESN* and ESN-TL predictions, respectively. In the relative-error metric, the ESN-TL prediction is more than twice as accurate as the ESN* control.
@@ 5 Discussion
We have shown that when the domain size $L$ and the dispersion parameter $\gamma$ are fixed, our ESN can learn and then accurately predict the power spectrum of the gKS equation. Further, we have shown that transfer learning allows one to efficiently and accurately track changes in the power spectrum of the gKS equation as $L$ or $\gamma$ varies. Importantly, our ESN did not blow up nor collapse during our long simulations. This is a great sign for the utility of ESNs with respect to long-term statistical prediction.
Looking to the future, ESNs can potentially be used as an inexpensive climate simulator or a chaotic heat bath. Importantly, ESNs allow relatively large timesteps when generating new trajectories, making them more efficient than traditional numerical integrators. Thus far, transfer learning for ESNs has received limited attention in the literature. We believe transfer learning is an important concept that can be applied to complex problems and allow ESNs to track parametric changes in spatiotemporally chaotic dynamics. In general, developing efficient models that can track parametric changes in dynamical systems is an important task. The methodology described in this paper aims at this goal.

### v2

@@ 4.1 Predicting the statistical behavior of the KS equation without transfer learning
Figure 2: Transfer learning allows the ESN to accurately predict the time-averaged power spectrum of the KS equation as the size of the spatial domain varies. Curves illustrate power spectra. DNS-L43: Power spectrum obtained by averaging over the $10$ DNS trajectories in the $L=43$ regime used for prediction. ESN-L22: Power spectrum for the $L=22$ regime predicted by the ESN, once it is trained on $20$ trajectories from the $L=22$ regime. ESN-TL: Power spectrum for the $L=43$ regime predicted by the ESN after transfer learning of level $10\%$ (left), $25\%$ (middle), and $50\%$ (right). Plots show $\log(e_{k})$ versus $\log(k)$, where $k$ denotes wavenumber.
In the absence of transfer learning, the ESN (trained only on trajectories from the $L=22$ regime) poorly predicts the power spectrum in the $L=43$ regime. However, transfer learning improves the prediction of the power spectrum in the $L=43$ regime. Figure 2 illustrates this improvement as a function of amount of TL data. The blue curve (same curve in each panel) is the power spectrum in the $L=43$ regime, computed by averaging over the $10$ DNS trajectories from this regime used for prediction. This curve serves as ground truth. The red curve (same curve in each panel) is the power spectrum for $L=22$ predicted by the ESN after having been trained on $20$ DNS trajectories from the $L=22$ regime. Comparing the red and blue curves, we see that the power spectrum for $L=22$ differs considerably from that for $L=43$. The green curves illustrate the ESN prediction of the power spectrum for the $L=43$ regime after transfer learning of level $10\%$ (left), $25\%$ (center), and $50\%$ (right). The quality of the prediction based on transfer learning rapidly improves as the level of TL data increases. In particular, the prediction is excellent at level $50\%$ (right).
We quantify the quality of the TL-based predictions of the power spectrum for $L=43$ by reporting the relative error
$\frac{|e_{k}^{\mathrm{DNS}}-e_{k}^{\mathrm{ESN}}|}{e_{k}^{\mathrm{DNS}}}$ (17)
associated with each of the first $15$ wavenumbers. See Table 6 and Figure 6 for this error data. Note the rapid improvement in the relative error associated with wavenumber $5$, the wavenumber that contains a plurality of the energy.
Transfer learning restores the ability of the ESN to predict structural features of the attractor as well. We consider one such feature in the paper, the correlation dimension of the attractor. Table 2 reports correlation dimension, computed using [5]. From left to right, the columns correspond to the red curve, then the three green curves, and finally the blue curve in Figure 2. Observe that the TL-based predictions of the correlation dimension approach ground truth (correlation dimension for $L=43$ computed using DNS data) as the amount of TL data increases.
Table 2: Transfer learning restores the ability of the ESN to predict the correlation dimension of the attractor when the length of the spatial domain instantaneously increases. From left to right, the columns correspond to the red curve, then the three green curves, and finally the blue curve in Figure 2.
ESN $L=22$ ESN TL $10\%$ ESN TL $25\%$ ESN TL $50\%$ DNS $L=43$
3.721 5.418 5.736 6.357 6.442
An ergodicity check. We performed a variant of the experiment described in Figure 2 in order to empirically verify ergodicity. This time, we changed the $20$ trajectories we used for training in the $L=22$ regime and we used one long DNS trajectory ($T=40000$) for prediction. Table 7 reports relative error as a function of wavenumber for this ergodicity experiment. As expected, relative errors in Tables 6 and 7 closely match for those wavenumbers that contain significant energy.
4.3 Transfer learning upgrades single-orbit prediction for the KS model
Transfer learning has been used with ESNs to successfully predict individual trajectories for shallow-water equations [13, 1]. In this section, we show that transfer learning meaningfully increases the length of time that single-orbit predictions for the KS model remain valid.
Suppose an ESN is trained on KS trajectories that correspond to spatial domain length $L=22$. How well would this ESN then predict single-orbit dynamics for larger values of $L$? Alam [1, Section 4.3.2] found that the timescale for which predictions of individual trajectories remain accurate decays approximately linearly with $L$ in this scenario. If $L>22$, one can potentially deal with this decay by training a fresh reservoir computing architecture on $L>22$ trajectories. Architecture options here include a larger ESN or a coupled network of ESNs. Transfer learning provides another potential solution. In this section, we show that transfer learning meaningfully enhances single-orbit predictive performance for the KS model when $L>22$.
We performed the following experiment. We initially trained the ESN using $20$ trajectories (each of length $T=20000$, or approximately $2000$ Lyapunov times) using $20$ KS trajectories from the $L=22$ regime. We used this particular ESN (no transfer learning) to then predict $5$ individual KS trajectories for integer increments of $L$ from $L=23$ to $L=29$. Table 3 (top) reports times (measured in Lyapunov times) for which relative prediction errors in $L^{2}$ remain below $10\%$. As expected, we observe decay as $L$ increases.
To assess the efficacy of transfer learning, we performed the following experiment for each integer value of $L$ from $L=23$ to $L=29$. We updated the training of the original ESN using one trajectory of length $20000$ from the new $L$ regime (a TL level of $5\%$ relative to the size of the original training dataset) and then used the new ESN to predict the same $5$ KS trajectories. Table 3 (bottom) reports times (measured in Lyapunov times) for which relative prediction errors in $L^{2}$ remain below $10\%$ for the TL-based predictions. Predictive performance typically improves by $1$ to $2$ Lyapunov times (bold indicates strict increase).
Figure 3 illustrates the difference between the ESN-predicted KS trajectory and the DNS-computed KS trajectory ($\mathrm{ESN}-\mathrm{DNS}$) without transfer learning (top row) and with transfer learning (bottom row) for Trajectory 2 in Table 3 with $L=25$ (left column) and $L=28$ (right column).
Table 3: Transfer learning improves single-orbit prediction for the KS model. We initially trained the ESN using trajectories from the $L=22$ regime. We used this ESN to predict $5$ trajectories for each integer value of $L$ from $L=23$ to $L=29$. The top grid reports Lyapunov times for which relative $L^{2}$-errors for these single-orbit predictions remain below $10\%$ (no transfer learning). After updating the training of the original ESN at a transfer learning level of $5\%$, ESN predictions typically remain accurate for 11 to 22 additional Lyapunov times. Bold indicates strict increase.
$L=23$ $L=24$ $L=25$ $L=26$ $L=27$ $L=28$ $L=29$
Trajectory 11 11 10 9 7 6 3 2
Trajectory 22 10 9 9 8 6 4 2
Trajectory 33 11 10 10 9 8 6 4
Trajectory 44 9 9 8 6 5 3 2
Trajectory 55 12 11 10 9 6 4 3
$L=23$ $L=24$ $L=25$ $L=26$ $L=27$ $L=28$ $L=29$
Trajectory 11 11 10 9 8 7 5 3
Trajectory 22 10 10 10 8 7 6 3
Trajectory 33 11 10 10 9 8 7 4
Trajectory 44 9 9 9 8 6 4 4
Trajectory 55 12 11 10 10 8 6 4
@@ 4.3 Transfer learning upgrades single-orbit prediction for the KS model
Figure 3: Difference between the ESN-predicted KS trajectory and the KS trajectory computed by direct numerical simulation without transfer learning (top row) and with transfer learning (bottom row). Left and right columns correspond to Trajectory 2 for $L=25$ and $L=28$, respectively, in Table 3.
4.4 Predicting the statistical behavior of the gKS equation
We consider the gKS equation when the parameter $\gamma$ in the dispersion relation is small. We assume that for $\gamma$ small, the gKS dynamics admit an attractor that supports an ergodic invariant measure. We assume that this invariant measure describes the asymptotic distribution of all of the initial data that we select. We show that for $\gamma$ small and fixed, the ESN accurately predicts the power spectrum of the gKS equation and the correlation dimension of the attractor for this fixed value of $\gamma$, once trained on trajectories generated using this fixed value of $\gamma$. We then show that transfer learning is remarkably effective for small $\gamma$. Throughout Section 4.4, we set $L=43$.
4.4.1 No transfer learning
We performed the following experiment for each of two values of $\gamma$, $\gamma=0$ and $\gamma=0.1$. We began with an untrained ESN. We generated $30$ gKS trajectories, each of length $T=10000$ (approximately $1000$ Lyapunov times), via direct numerical simulation. We used $20$ of these to train the ESN and the other $10$ for prediction.
Figure 4 illustrates that for $\gamma=0$ (left) and $\gamma=0.1$ (right), the trained ESN accurately predicts the power spectrum of the gKS equation. Each red curve shows the ESN prediction of the power spectrum. Each blue curve shows the power spectrum obtained by averaging over the $10$ DNS trajectories used for prediction. Table 4 reports that the trained ESN accurately captures the correlation dimension of the attractor as well. Importantly, the power spectrum of the gKS equation and the correlation dimension of the attractor depend sensitively on $\gamma$ in the chaotic (small $\gamma$) regime. This observation motivates the use of transfer learning in this regime.
Refer to caption Refer to caption
@@ 4.4.1 No transfer learning
Figure 4: Once trained, the ESN accurately predicts the time-averaged power spectrum of the gKS equation. Left: $\gamma=0$. Right: $\gamma=0.1$. Each blue curve shows the power spectrum obtained by averaging over the $10$ DNS trajectories (each of length $T=10000$) used for prediction. Each red curve shows the power spectrum predicted by the ESN, after the untrained ESN has been trained using $20$ trajectories (each of length $T=10000$). Plots show $\log(e_{k})$ versus $\log(k)$, where $k$ denotes wavenumber. Spatial domain: $L=43$.
Table 4: Once trained, the ESN accurately predicts the correlation dimension of the attractor of the gKS equation. These correlation dimensions correspond to Figure 4.
$\gamma=0$ $\gamma=0.1$
DNS 6.442 3.416
ESN 6.311 3.449
4.4.2 Using transfer learning for the gKS equation
Figure 4 demonstrates that the statistical behavior of the gKS equation depends sensitively on $\gamma$ in the small-$\gamma$ regime. We therefore ask if transfer learning can capture substantial changes in statistical behavior as $\gamma$ varies. To address this question, we performed the following experiment. We started with an untrained ESN and first trained it using $20$ trajectories (each of length $T=10000$) from the $L=43$ and $\gamma=0$ regime. We then updated the training using some TL data from the $\gamma=0.1$ regime at transfer learning rate $\alpha=0.005$. After the training update, we predicted the power spectrum of the gKS equation for $\gamma=0.1$.
@@ 4.4.2 Using transfer learning for the gKS equation
The result of this experiment is shown in Figure 5. The blue curve shows the ground truth, namely the power spectrum in the $\gamma=0.1$ regime obtained by averaging over the $10$ DNS trajectories (each of length $T=10000$) from the $\gamma=0.1$ regime used for prediction. Using a transfer learning level of only $10\%$ (22 trajectories) produces an accurate ESN prediction of the power spectrum (green curve). For reference, the yellow curve shows the ESN-based prediction of the power spectrum when the ESN is trained only on $20$ trajectories from the $\gamma=0$ regime (no transfer learning). See Table 8 and Figure 7 for relative error (17) as a function of wavenumber. Importantly, the TL-enhanced ESN can continue to accurately predict the power spectrum over long timescales. By contrast, single-orbit predictions in the $\gamma=0.1$ regime based on transfer learning are valid for much shorter timescales [1, Section 5.1.1].
@@
As a control, we compare the prediction using transfer learning (Figure 5, green curve, ESN-TL) to a prediction from an ESN that has been trained only on the TL data, namely $2$ trajectories from the $\gamma=0.1$ regime (Figure 5, red curve, ESN*). Visually (Figure 5) and quantitatively (Table 8 and Figure 7), the ESN* prediction is inferior to the ESN-TL prediction. Discrepancies between these two predictions occur at low wavenumbers that contain most of the energy. The curves shown in Figure 5 contain total energies $2.92$ (blue, DNS), $2.80$ (red, ESN*), and $2.97$ (green, ESN-TL). This corresponds to relative error $4.1\%$ and $1.7\%$ for the ESN* and ESN-TL predictions, respectively. In this sense, the ESN-TL prediction is more than twice as accurate as the ESN* control.
Table 5 reports correlation dimension of the attractor for the experiment from Figure 5. We see that transfer learning allows the ESN to accurately predict correlation dimension for $\gamma=0.1$, and the ESN-TL prediction is better than the control (ESN*) prediction.
Figure 5: Transfer learning allows the ESN to accurately predict the time-averaged power spectrum of the gKS equation as the dispersion parameter $\gamma$ instantaneously jumps. Blue curve: Power spectrum of the gKS equation in the $\gamma=0.1$ regime obtained by averaging over the $10$ DNS trajectories (each of length $T=10000$) used for prediction. Green curve: Power spectrum predicted by the ESN after a first round of training using $20$ trajectories from the $\gamma=0$ regime (each of length $T=10000$) and then a training update using TL data from the $\gamma=0.1$ regime at level $10\%$. Yellow curve: Power spectrum predicted by the ESN after training only on $20$ trajectories from the $\gamma=0$ regime (no transfer learning). Red curve: Power spectrum predicted by an ESN that has been trained only on the TL data, namely $2$ trajectories from the $\gamma=0.1$ regime. Plots show $\log(e_{k})$ versus $\log(k)$, where $k$ denotes wavenumber. Spatial domain: $L=43$.
Table 5: Correlation dimension of the attractor of the gKS equation for the experiment from Figure 5.
DNS ($\gamma=0.1$) ESN ($\gamma=0$) ESN* ($\gamma=0.1$) ESN-TL
3.416 6.311 3.242 3.401
@@ 5 Discussion
We have shown that when the domain size $L$ and the dispersion parameter $\gamma$ are fixed, our ESN can learn and then accurately predict the power spectrum of the gKS equation and the correlation dimension of the attractor. Further, we have shown that transfer learning allows one to efficiently and accurately track changes in the power spectrum of the gKS equation and the correlation dimension of the attractor when $L$ or $\gamma$ instantaneously jumps. Importantly, our ESN did not blow up nor collapse during our long simulations. This is a great sign for the utility of ESNs with respect to long-term statistical prediction.
Our results suggest that transfer learning is more efficient when the correlation dimension of the attractor drops upon a parameter set jump than when it rises upon such a jump. In Section 4.2, we initially train the ESN using KS trajectories for $L=22$ and then apply transfer learning to do statistical prediction for the KS equation when $L=43$. The jump from $L=22$ to $L=43$ induces a jump in the correlation dimension of the attractor (Table 2). In this scenario, we observe that a transfer learning level of $50\%$ produces accurate statistical predictions. By contrast, in Section 4.4.2 we fix $L=43$, initially train when $\gamma=0$, and then apply transfer learning to do statistical prediction for the gKS equation when $\gamma=0.1$. This time, the jump from $\gamma=0$ to $\gamma=0.1$ induces a drop in the correlation dimension of the attractor (Table 5). In this second scenario, we observe that a transfer learning level of only $10\%$ suffices for accurate statistical predictions. It is intuitive that transfer learning is more efficient when the correlation dimension of the attractor decreases upon a parameter set jump than when it rises upon such a jump, since the output layer $W_{\mathrm{out}}$ must be updated to project higher-dimensional chaos in the latter case.
Looking to the future, ESNs can potentially be used to inexpensively simulate climate or to generate chaotic timeseries with known statistical properties. Importantly, ESNs allow relatively large timesteps when generating new trajectories, making ESNs more efficient than traditional numerical integrators. Thus far, transfer learning for ESNs has received limited attention in the literature. We believe transfer learning is an important concept that can be applied to complex problems and allow ESNs to track parametric changes in spatiotemporally chaotic dynamics. In general, developing efficient models that can track parametric changes in dynamical systems is an important task. The methodology described in this paper aims at developing efficient computational ML models for this task.
@@ 7 Author contributions
8 Supporting tables and figures
Table 6: Relative error (17) as a function of wavenumber for the TL-based predictions in Figure 2.
Wavenumber ESN-L22 ESN-TL 10% ESN-TL 25% ESN-TL 50%
1 0.603 0.445 0.194 0.048
2 10.496 0.771 0.307 0.131
3 4.403 1.022 0.248 0.014
4 0.468 0.336 0.101 0.034
5 0.943 0.787 0.381 0.004
6 0.965 0.495 0.118 0.004
7 0.989 0.536 0.211 0.128
8 0.997 0.319 0.072 0.113
9 0.999 0.308 0.237 0.047
10 1.000 0.465 0.137 0.026
11 1.000 0.416 0.031 0.013
12 1.000 0.303 0.197 0.054
13 1.000 0.324 0.111 0.062
14 1.000 0.298 0.064 0.072
15 1.000 0.585 0.132 0.035
Figure 6: Plot corresponding to Table 6.
Table 7: Relative error (17) as a function of wavenumber for the ergodicity experiment described in Section 4.2.
Wavenumber ESN-L22 ESN-TL 10% ESN-TL 25% ESN-TL 50%
1 0.611 0.432 0.193 0.046
2 10.315 0.770 0.313 0.136
3 4.372 1.005 0.249 0.014
4 0.459 0.348 0.102 0.033
5 0.927 0.766 0.377 0.005
6 0.971 0.480 0.122 0.013
7 0.982 0.518 0.211 0.125
8 0.998 0.316 0.073 0.112
9 1.000 0.316 0.242 0.047
10 1.000 0.463 0.140 0.025
11 1.000 0.429 0.031 0.013
12 1.000 0.294 0.198 0.055
13 1.000 0.325 0.110 0.062
14 1.000 0.305 0.064 0.071
15 1.000 0.590 0.135 0.036
Table 8: Relative error (17) as a function of wavenumber for the experiment in Figure 5.
Wavenumber ESN $(\gamma=0)$ $^{*}(\gamma=0.1) ESN-$TL
1 0.232 0.137 0.036
2 0.699 0.178 0.07
3 0.742 0.161 0.048
4 0.253 0.118 0.081
5 1.702 0.187 0.028
6 0.915 0.470 0.036
7 0.048 0.025 0.028
8 0.808 0.023 0.027
9 0.077 0.021 0.021
10 1.097 0.050 0.083
11 0.188 0.128 0.006
12 0.795 0.192 0.08
13 0.676 0.025 0.012
14 0.267 0.051 0.009
15 0.452 0.167 0.031
Figure 7: Plot corresponding to Table 8.